OpenStudy (anonymous):

Which is a requirement for a number to a rational number? The number can be written as a fraction. The denominator is an integer. The numerator is an integer. All of the above.

OpenStudy (phi):

\[ \frac{1}{0.5} \] is rational (equal to 2) similarly \[ \frac{0.5}{1}\] is rational = 1/2

OpenStudy (anonymous):

@ImNotGoodAtMath

jimthompson5910 (jim_thompson5910):

Let's go through all these choices A) "The number can be written as a fraction." A fraction of what? If we have a fraction of numbers like \(\large \sqrt{2}\) and \(\large \sqrt{7}\), then we don't have a rational number. The number has to be able to be written as a fraction of INTEGERS This all shows that choice A is false. ------------------------------------------------------- B) " The denominator is an integer." This is a slightly better clarification. So the denominator is now an integer, but what about the numerator? That could be a number like \(\large \sqrt{31}\), which makes the number not rational. This shows B is wrong. ------------------------------------------------------- C) " The numerator is an integer." All integers are rational numbers because you can use the rule x = x/1 Examples 3 = 3/1 41 = 41/1 88 = 88/1 -23 = -23/1 etc etc So because we can write any integer as a fraction of two integers, this means that all integers are rational numbers. It doesn't work the other way around since 1/2 is rational but not an integer. So we have our answer here. ------------------------------------------------------- D) " All of the above." This is automatically false once you've proven A is false

OpenStudy (anonymous):

ty i was lost and im very thankfull you took ur time to explian every answer :)